OpenStudy (anonymous):

using the Mean Value Theorem, find the integral of square root of x with the values of [4,9], i have the formula but, i am gettin lost with the formula

OpenStudy (anonymous):

\[F \prime = \frac{ f(b)-Fa) }{ b-a }\] just substitute and their conditions is met with this interval

OpenStudy (anonymous):

b = 9 a = 4 but are you sure the question asked you about the integration??

OpenStudy (anonymous):

sorry, umm the formula is \[\int\limits_{a}^{b}f(x)dx=f(c)[b-a]\]

OpenStudy (anonymous):

\[5\sqrt{x}\]

OpenStudy (anonymous):

To clarify... you want to evaluate the integral \[\int_4^9\sqrt x~dx\] using the MVT, which says if a function is continuous in the closed interval [a,b] and differentiable in the open interval (a,b), then there is some c between a and b such that \[f'(c)=\frac{f(b)-f(a)}{b-a}\] In this case (for the integral MVT), \[f(b)-f(a)=\int_a^b f'(x)~dx~~\iff~~f(9)-f(4)=\int_4^9\sqrt x~dx\]

OpenStudy (anonymous):

but that isn't the MVT at all dx doesn't equal 5 !!! MVT is the same as linear approximation except it gives an exact value .

OpenStudy (anonymous):

The MVT you're familiar with (involving derivatives) is the same for integrals with a slight variation. I'm not sure what you're trying to say @Catch.me

OpenStudy (anonymous):

yeah he is right

OpenStudy (anonymous):

what i meant that\[dx \neq b-a =\] look at the last formula in that link http://www.cut-the-knot.org/Curriculum/Calculus/MVT.shtml

OpenStudy (anonymous):

No one's claiming that \(dx=b-a\). The question itself is kinda weird to begin with, @nutterbutter. The MVT says \[F'(c)=\frac{F(b)-F(a)}{b-a}~~\Rightarrow~~f(c)=\frac{1}{b-a}\int_a^b f(t)~dt\] In order to find the integral, you need the average value, but to find that you need to know the integral... See the problem here?

OpenStudy (anonymous):

OK and what he asked is to find the integration by MVT,but you wrote \[\int\limits_{a}^{b}\sqrt{x} dx =F(9) - F(4)\] which isn't due to MVT but what related is that \[\int\limits_{a}^{b}\sqrt{x} dx = f(c)(b-a).= 5 \sqrt{x}.\]

OpenStudy (anonymous):

or I got you wrong I am sorry. I think the answer is 5 sqrt x

OpenStudy (anonymous):

i meant what sithsandgiggles is saying

OpenStudy (anonymous):

The answer must be a number. A definite integral is always a number.

OpenStudy (anonymous):

OK then you will integrate \[\sqrt{X}\] to get F(b) - F(a) right??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thus you didn't use MVT can you understand me ?? MVT didn't proved F(b)-F(a) = integral f(x). ,so your methods won't use MVT and that is my point MVT said integral/b-a = f(c) -----> integral ={b-a}f(c).

OpenStudy (anonymous):

.........

OpenStudy (anonymous):

you didn't get me yet ??

OpenStudy (anonymous):

This is what I mean MVT says about the integration that: \[\int\limits_{a}^{b} f(x)dx = \left[ b-a \right] f(c).\]